please help me on this one.. if 2t + 3 = -9 what is the value of -3t - 7.....theres one and the next ? is solve 2x - 4 = 7 justify each step

Answer by pwac(253)

First find t
2t+3=-9 Subtract 3 from both sides
2t=-12 Divide both sides by 2
t=-6
Now put this value of t into -3t-7
-3(-6)=18
18-7=11
2x-4=7 Add 4 to both sides:
2x=11 Divide both sides by 2
x=11/2=5.5
Pete